OpenStudy (anonymous):

Interest Problem: Julian deposits money into two different savings accounts. He deposits $600 more into Account B than Account A. After a certain amount of time, Account A has earned $180 at a rate of 4%, and Account B has earned $375 at a rate of 5%. How much did Julian initially deposit into each account? What equation is used to solve this problem? What does each variable represent? Identify the given information. What do you know about P? Enter the given information into the equation. Solve the equation. State the solution.

OpenStudy (anonymous):

in I over pr form. I=interest earned, P=initial amount deposited R=interest rate.

OpenStudy (slaaibak):

Say P is the initial amount account A P + 600 is the initial amount in B First equation: P * .04 * t = 180 Second equation: (P + 600) * 0.05 * t = 375 so just solve simultaneously for t and P

OpenStudy (anonymous):

but that isn't in fraction form...

OpenStudy (slaaibak):

i'm giving you guidelines on how to do this, not going to do the whole thing for you.

OpenStudy (anonymous):

well that isn't really helping me... i'm not understanding how to do it with it in the fraction. what i have is 180 over P(.04) = 375 over (P+600)(.05)

OpenStudy (slaaibak):

that's correct. \[{ 180 \over 0.04P} = {375 \over (P + 600)(.05)} \rightarrow 180(0.05)(P + 600) = (375)(0.04)P\] just solve that equation to get P

OpenStudy (anonymous):

So, 900 dollars is the initial amount deposited into account A, and 1500 into account B?

OpenStudy (anonymous):

???

OpenStudy (slaaibak):

yes

OpenStudy (anonymous):

Thank you so much (: i'll give you best response.

OpenStudy (slaaibak):

no problem